The assets of Dallas & Associates consist entirely of current assets and net plant and equipment, and the firm has no excess cash. The firm has total assets of $2.6 million and net plant and equipment equals $2.2 million. It has notes payable of $155,000, long-term debt of $754,000, and total common equity of $1.55 million. The firm does have accounts payable and accruals on its balance sheet. The firm only finances with debt and common equity, so it has no preferred stock on its balance sheet.

Write out your answers completely. For example, 25 million should be entered as 25,000,000. Negative values, if any, should be indicated by a minus sign. Round your answers to the nearest dollar, if necessary.

What is the company's total debt?

What is the amount of total liabilities and equity that appears on the firms balance sheet?

What is the balance of current assets on the firm's balance sheet?

What is the balance of current liabilities on the firm's balance sheet?

What is the amount of accounts payable and accruals on its balance sheet? (Hint: Consider this as a single line item on the firm's balance sheet.)

What is the firm's net working capital? If your answer is zero, enter "0". Negative value, if any, should be indicated by a minus sign.

What is the firm's net operating working capital?

What is the monetary difference between your answers to part f and g?

What does this difference indicate?

a. notes payable balance

b. accounts payable balance

c. current liabilities balance
Answer #1

a.Total Debt = Total Assets – Total Equity

= 2,600,000 – 1,550,000

= $1,050,000

b.Total liabilities and Equity is equal to total assets

= $2,600,000

c.Current Assets = Total Assets – Plant and Equipment

= 2,600,000-2,200,000

= 400,000

d.Current Liabilities = Total Liabilities – Long term debt

= 1,050,000 – 754,000

= $296,000

e.Accounts payables and accruals = current liabilities – notes payables

= 296,000 – 155,000

= $141,000

f.Working capital = Current Assets – Current Liabilities

= 400,000-296,000

=104,000

Net operating working capital = current assets – Accounts payables and accruals

= 400,000 – 141,000

= 259,000

h.Difference = 104,000-259,000

= -155,000

Difference indicates note payable balance